### v2

# ANote[fnsymbol]

Abstract
Human beings perceive the real world through a spectrum of sensory modalities, encompassing auditory, visual, and linguistic faculties. The journey towards achieving Artificial General Intelligence (AGI) necessitates the development of models that can emulate these multifaceted perceptual capabilities and comprehensively understand these diversified data. To this end, we introduce Nexus-O, an industry-level omni-perceptive and -interactive model capable of efficiently processing Audio, Image, Video, and Text data in any combination and output audio/text in an end-to-end way. We systematically investigate Nexus-O by addressing three key research questions: First, how can models be efficiently designed and trained to achieve tri-modal alignment, understanding and reasoning capabilities across multiple modalities? Second, what approaches can be implemented to evaluate tri-modal model robustness, ensuring reliable performance and applicability in real-world scenarios? Third, what strategies can be employed to curate and obtain high-quality, real-life scenario speech datasets? For the first question, we design and pre-train Nexus-O based on the vision-language model, rather than the language model. By pre-training the model over high-quality synthetic audio data, our model is capable of tri-modal perception and interaction. For the second question, we introduce a new audio testbed, Nexus-O-audio, comprising diverse Automatic Speech Recognition (ASR) samples, spanning various real-world scenarios, such as corporate meetings and live stream. For the third question, we design the speech data synthesis pipeline to obtain high-quality speech training datasets, covering various real-world scenarios. Comprehensive experimentation and an in-depth analysis of tri-modal alignment over latent space demonstrate the advantages of our model on downstream tasks.
11footnotetext: * denotes equal contribution. This paper was completed during Che Liu and Shilin Zhou’s internships at Hithink Research. 11footnotetext: $\spadesuit$ denotes Project Leader. $\dagger$ denotes corresponding author, daiyongya@outlook.com.
1 Introduction
Figure 1: An overview of the model’s performance. Nexus-O is capable of handling a wide range of tasks, demonstrating strong performance across various benchmarks. For Nexus-O-audio (EN) and (CN), the scores are expressed as reciprocal (1/scores).
Since the mid-20th century, significant efforts have been devoted to designing systems of the capability to simulate human cognition, with the ultimate goal of developing Artificial General Intelligence (AGI) (Feng et al., 2024; Morris et al., 2023). Notable algorithms include Rule-based systems, Neuro-symbolic AI systems (Sheth et al., 2023), Bayesian nets systems (Heckerman, 2022), Neural networks systems (Hinton, 2006), etc. Among them, neural networks systems have emerged as a cornerstone thanks to their remarkable generalization capabilities. In this context, various neural architectures have been proposed, such as CNN (LeCun et al., 1995) and Transformer (Vaswani, 2017). The discovery of neural scaling laws (Kaplan et al., 2020) has further advanced the field, paving the way for Transformer-based Large Language Models (LLMs), such as ChatGPT (Kaplan et al., 2020), which centre on language as the medium for perceiving and interacting with both human society and the physical world. By aligning multiple modalities around the language, Multimodal Large Language Models (MLLMs) have gained widespread recognition as a promising pathway toward AGI (Liang et al., 2024; Zhang et al., 2024a; NVIDIA et al., 2025; Ma et al., 2025; Hong et al., 2025).
MLLMs, involving language, vision, and auditory modalities, can generally be divided into three categories: vision-language model (Wang et al., 2024), audio-language model (Chu et al., 2024), and vision-language-audio model (Fu et al., 2025). While alignment among three modalities offers a more comprehensive simulation of human cognitive processes than two-modality alignment, it presents significantly greater challenges. These challenges include limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. These limitations hinder the advancement of omni-perceptive and interactive AGI systems. To address these challenges, we explores the development of a fully end-to-end tri-modal system capable of omni-perception and seamless interaction with the world in this work.
To achieve this goal, several key research questions need to be addressed. First, what methodologies and strategies can be employed to curate and obtain high-quality, real-scenario speech datasets? Second, how can models be efficiently designed and trained to achieve tri-modal alignment, understanding and reasoning capabilities across multiple modalities? Third, what approaches can be implemented to evaluate audio-modal robustness, ensuring reliable performance and applicability in real-world scenarios?
In response to these questions, we propose Nexus-O with the following three approaches: (1) We systematically design the speech data synthesis pipeline to obtain high-quality speech training datasets, spanning various real-world scenarios, such as corporate meetings, live broadcasting scenarios, etc. (2) We start by designing and pre-training Nexus-O based on the vision-language model, especially Qwen2.5-VL (Bai et al., 2025) [^1] In this report, Qwen2.5-VL refers to the instruction version., rather than a language model, as is typically chosen in most omni models (Fu et al., 2024c; Xie & Wu, 2024a; b; Li et al., 2024a). By pre-training the model over high-quality synthetic audio data, Nexus-O is capable of tri-modal perception and interaction. This strategy offers high training efficiency while reducing the computational resource burden derived from the training vision-language alignment stage, which is an efficient and effective option for most institutions with low computational resources. (3) For a fair evaluation of the model in real-world scenarios, we introduce a new audio testbed, i.e., Nexus-O-audio, comprising diverse Automatic Speech Recognition (ASR) samples, spanning various real-world scenarios.
The experimental results demonstrate the following five key findings: (1) In the visual understanding task, Nexus-O exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, indicating the effectiveness of our training strategy. (2) Within the Audio English QA benchmark, the model achieves better accuracy than the same-period competitor (i.e, MiniCPM-o2.6-7B) in the LLaMA Q. benchmark (Nachmani et al., 2024). (3) In our real-world ASR benchmarks, Nexus-O achieves outstanding performance, which underscores its robustness in real scenarios. (4) In the Speech-to-Text Translation benchmark, our model outperforms Qwen2-Audio-Instruct-7B. (5) Furthermore, an in-depth analysis of tri-modal alignment reveals that incorporating the audio modality enhances representational alignment. Even in the absence of audio input, our model delivers performance comparable to the backbone model. The main contributions of this paper are summarized as follows:
  • We systematically design the speech data synthesis pipeline to obtain high-quality speech datasets to facilitate the rapid deployment of the model in real-world scenarios.
  • We propose an omni-perceive and -interactive tri-modal model, i.e., Nexus-O, which supports any combination of audio, image/video, and text inputs. Different from existing approaches that treat audio as an auxiliary, our model enables a joint understanding of these modalities.
  • To evaluate the performance of the tri-modal model, we propose the Nexus-O-audio testbed, covering various real-world scenarios, such as corporate meetings, live broadcasting scenarios, etc. We intend to release the Nexus-O-audio testbed to establish a new standard for fostering innovation in addressing complex real-world applications.
  • We comprehensively evaluate the performance of Nexus-O on various benchmarks and conduct an in-depth analysis of tri-modal alignment from the perspective of representation space. Experimental results illustrate the efficiency and the effectiveness of our tri-modal model.
The structure of this work is organised as follows: In Section 2, we first identify the research gap between our model and other contemporaneous approaches, and then discuss the limitations of current speech benchmarks. In Section 3, we present our tri-modal model architecture, the composition of training data and audio data synthesis pipeline, pre-training strategies, and the Nexus-O-audio testbed. In Section 4, we first address the alignment of downstream tasks, including vision-language evaluation and audio-language evaluation, and then evaluate representational alignment. Finally, we conclude with a summary of our findings and outline our future research directions in Section 5.
2 Related Work
In this section, we present the existing works of Multimodal Large Language Models (MLLMs) in Section 2.1, and analyse existing audio modality evaluation approaches in Section 2.2.
2.1 MLLMs
Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and reasoning over textual knowledge (Touvron et al., 2023a; b; Dubey et al., 2024; Yang et al., 2024a; Jiang et al., 2023). Building on these advancements, a bunch of studies (Chen et al., 2024a; b; Dai et al., 2022) have extended the understanding and alignment capabilities of LLMs to visual knowledge, addressing challenges such as vision-language alignment and instruction-following, by introducing Multimodal Large Language Models (MLLMs). Notable models include Qwen-VL (Wang et al., 2024; Yang et al., 2024a), InternVL (Chen et al., 2024c), etc. However, in complex visual reasoning tasks, such as video analysis, MLLMs typically struggle to perform well due to the heterogeneity among diverse modalities, where different modalities have distinct feature granularities, representations, and geometries in a shared latent space (Zhang et al., 2024e; c). To further improve their performance in visual understanding tasks, recent works have expanded MLLMs by incorporating audio capabilities, such as LLaMA-Omni (Fang et al., 2024a), VITA-1.0 (Fu et al., 2024c), Mini-Omni2 (Xie & Wu, 2024b), Baichuan-Omni (Li et al., 2024a), and MinMo (Chen et al., 2025a). However, these models exhibit certain limitations: LLaMA-Omni and MinMo lacks visual capabilities, Baichuan-Omni does not have end-to-end Text To Speech (TTS) capabilities, and VITA-1.0 is constrained by the limited capability of its backbone model.
Recent models, such as VITA-1.5 (Fu et al., 2025), MiniCPM-o2.6  [^2] https://github.com/OpenBMB/MiniCPM-o, Baichuan-omni-1.5333https://github.com/baichuan-inc/Baichuan-Omni-1.5, and OpenOmni (Luo et al., 2025), have been proposed and have achieved state-of-the-art performance. However, these are pretrained based on the LLMs rather than MLLMs. While pre-training from scratch can enhance multimodal alignment, this strategy significantly increases training time and resource demands due to the necessity of additional visual-language alignment, which relies on large-scale pre-training datasets. To improve training efficiency and to reduce resource burden, we start by pre-training over MLLM, specifically Qwen2.5-VL-7B, and introduce Nexus-O, an advanced unified tri-modal LLM designed to enable audio understanding and generation across arbitrary combinations of modalities. This strategy is friendly with institutions with limited computation resources.
2.2 Audio Modality Evaluation
With the rapid advancement of MLLMs, a wide range of multimodal benchmarks has been proposed and widely adopted for evaluation, ranging over vision (Xie et al., 2023; Gan et al., 2024; Fu et al., 2024b; Yue et al., 2024; Yang et al., 2025), audio (Conneau et al., 2023; Yang et al., 2024b; Fu et al., 2024a), omni(Li et al., 2024b), etc. In the speech domain, existing benchmarks, e.g., Fleurs (Conneau et al., 2023), Aishell2 (Du et al., 2018a), LibriSpeech (Pratap et al., 2020b), and Common Voice (Ardila et al., 2019), provide linguistically diverse datasets under varying acoustic conditions, facilitating Automatic Speech Recognition (ASR) evaluation. Additionally, AIR-Bench (Yang et al., 2024b) introduces a pioneering generative evaluation framework encompassing speech, natural sounds, and music, employing novel audio mixing strategies and GPT-4 for unified, objective, and reproducible assessments.
However, the assessment of speech models should be closely aligned with real-world environments to ensure that the evaluation results faithfully represent the demands and requirements of practical and industrial applications. We observe that these existing speech benchmarks present three limitations. First, the small data scale of these benchmarks typically results in significant performance variance. Second, the benchmarks primarily focus on controlled scenarios, making it difficult to evaluate models in dynamic and complex real-world settings, e.g., meetings or live broadcast environments. Third, scores from the existing ASR benchmarks generally fail to align closely with the actual user interaction experience, thereby limiting their practical utility. To address these limitations, we propose Nexus-O-audio testbed, comprising 8.1k and 2.8k of Chinese and English ASR samples, spanning various application domains, such as corporate meetings and live broadcasting scenarios.
In addition to assessing audio downstream task alignment, we conduct an in-depth evaluation of whether audio modality can enhance vision-language alignment, as well. Intuitively, the audio modality can provide complementary information to the visual and language modalities, enhancing mutual information and feature alignment, thereby improving the performance on visual understanding tasks (Song et al., 2023; Choi et al., 2024; Zhang et al., 2024b).
Figure 2: Architecture of Nexus-O, which is designed to accept any combination of input modalities and generates output in either the language or audio modality, where the Auto-Regressive (AR) audio decoder takes a special start token embedding and the last language embedding as input to generate the hierarchical discrete audio codes in an auto-regressive manner. These codes are subsequently fed into a pretrained audio generator to produce the final audio output.
3 Nexus-O
In this section, we introduce the architecture of Nexus-O in Section 3.1, training data composition and audio data synthesis pipeline in Section 3.2, pre-training strategies in Section 3.3, and Nexus-O-audio testbed in Section 3.4, respectively.
3.1 Architecture
Visual encoder.
Following Qwen2.5-VL (Bai et al., 2025), the vision encoder employs a re-designed Vision Transformer (ViT) architecture by incorporating M-RoPE and window attention to support native input resolutions while accelerating the computation of the entire visual encoder. During both training and inference, the height and width of the input images are resized to multiples of 28 before being fed into the ViT. The vision encoder processes images by splitting them into patches with a stride of 14, generating a set of image features.
Audio encoder/decoder.
To enable audio capabilities in the vision-language MLLM, we incorporate an audio encoder-decoder architecture. In this setup, the encoder is responsible for mapping the speech features into the semantic space of the MLLM, while the decoder transforms the semantic code back into speech, as shown in Figure 2. Specifically, the audio encoder comprises a pre-trained Whisper-large-v3 (Radford et al., 2023), and a two-layer MLP adapter. Specifically, the first layer of the MLP concatenates five consecutive speech feature vectors into a single vector, which is subsequently processed by the second MLP layer.
The auto-regressive audio decoder is a 6-layer decoder-only transformer designed to generate discrete speech codes, which are then passed to a pre-trained generator to produce the final waveforms. Specifically, we use the audio codebook from the FireFly GAN in Fish-Speech (Liao et al., 2024) as the audio codebook. Following MusicGen (Copet et al., 2023), the audio decoder predicts sequences of eight speech codes in a delayed manner. At each time step, the decoder averages the embeddings of the eight speech codes generated in the previous time step with the hidden state of the LLM at the current time step. The resulting averaged embedding is then fed into the audio transformer to predict the next eight speech tokens. Finally, the predicted speech tokens are input into the generator of the FireFly GAN to synthesise the final waveforms.
Language model.
Nexus-O utilises the language model component of Qwen2.5-VL-7B (Bai et al., 2025) as the foundational language model, consisting of 28 causal Transformer layers.
3.2 Training Data
Figure 3: Audio dataset synthesis pipeline. In the current version, our testbed only supports the ASR task. We will further incorporate various audio-modal tasks, including the AQA and AVQA tasks. Both are currently under development.
Datasets composition.
As detailed in Table 1, the training data for Nexus-O encompasses a wide range of domains within speech processing. These domains include recognition tasks such as ASR and Speech Emotion Recognition (SER), reasoning tasks include Automatic Audio Caption and Speech Emotion Recognition, translation tasks like Speech-to-Text Translation (S2TT), generation tasks such as Speech Synthesis, and interactive tasks involving Speech-to-Speech Chat and Speech-to-Text Chat. The pretrained datasets include (1) public corpora (50%): Aishell1 (Bu et al., 2017), Wenetspeech (Zhang et al., 2022), Librispeech (Panayotov et al., 2015a), Kensho (O’Neill et al., 2021), PeopleSpeech (Galvez et al., 2021), MLS (Pratap et al., 2020a), and Gigast (Ye et al., 2023) and (2) real-life in-house data (50%).
Table 1: Datasets description. The target model for each stage is described in Figure 4. The template for each task is provided in Appendix A.
Stage Task Description Num
1 Automatic Speech Recognition 5.6M
Speech-to-Text Translation 1.5M
Speech-to-Text Chat 0.2M
Automatic Audio Caption 0.14M
Speech Emotion Recognition 0.75M
2 Speech-to-Text Chat 0.23M
Automatic Speech Recognition 20k
Speech-to-Text Translation 20k
3 Speech Synthesis 40M
Speech-to-Speech chat 0.22M
Audio dataset synthesis pipeline.
For the non-speech datasets, we employ the zero-shot in-context generation technique proposed by CosyVoice (Du et al., 2024) to convert the text into natural-sounding speech. To enhance the diversity of voices, we randomly select a speaker from a pool of two thousand individuals for each corresponding language to synthesise the speech, ensuring that the voice aligns with the language of the text. However, many conversational and instruction datasets are not ideal for speech synthesis due to the nature of the tasks, which do not facilitate effective speech interaction. To address this issue, we enhance the quality of our synthesised speech data by implementing the following optimisation measures:
  • Length Filtering: texts exceeding a predefined length threshold are excluded from the dataset. Specifically, texts longer than 200 Chinese characters or 200 English words are removed.
  • Non-text Element Filtering: texts containing an excessively high proportion of non-text elements are excluded. Specifically, texts with a non-text character ratio exceeding 0.2 are removed. Text characters in this context include punctuation marks (e.g., !, ?), alphabetic characters ([a-z]), and Chinese characters.
  • Pattern Matching Filtering: texts that included URLs, file paths, or common LaTeX formulas were identified and removed using regular expressions.
We further enhance the data quality with the help of a large model through the following steps:
  • Filter out questions that are not suitable for oral QA.
  • Rewrite the command in a conversational tone and generate a conversational response.
3.3 Training Strategies
Figure 4: Overview of the training stage in Nexus-O. The first stage aims to map the speech features into semantic space, the second stage aims to enable the audio instruction-following capability, and the last stage aims to enable speech generation capability.
Stage 1: audio alignment.
The first stage aims to facilitate the capacity of LLMs to comprehend and process input speech features. To accomplish this, we utilise 8 million bilingual speech-text pairs for alignment training, covering five training objectives, as illustrated in Table 1. Consistent with the methodology employed in Qwen2.5-VL-7B, our training data is structured in the ChatML format. An example of this format is illustrated as below:
The Dataset Format Example of ChatML <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|>Please help me transcribe my current speech into text. <|im_end|> <|im_start|>assistant hat had its source away back in the woods of the old cuthbert place it was reputed to be an intricate headlong brook in its earlier course through those woods with dark secrets of pool and cascade but by the time it reached lynde’s hollow it was a quiet well conducted little stream <|im_end|>
Stage 2: audio SFT.
The second stage aims to enable the model to follow speech instructions by incorporating speech-based questions and text-based answers. In this stage, we freeze the visual and audio encoders while unfreezing the visual and audio adapters, and backbone LLM. This strategy is designed to enhance the capability of Nexus-O to respond accurately to multimodal instructions, thereby improving its performance in understanding and processing both visual and auditory information.
By keeping the core encoder parameters fixed, we ensure that the foundational representations remain stable while allowing the adapters to fine-tune and specialize in the nuances of specific tasks or instructions. This strategic adjustment aims to refine the sensitivity of Nexus-O to multimodal cues without disrupting the robustness of its underlying architecture.
Stage 3: audio output.
Different from the previous large speech model connected to the TTS module (Fu et al., 2024c; Li et al., 2024a), we exploit a decoder to generate speech end-to-end. First, we train Nexus-O on the Text-to-Speech dataset to make the model learn text-to-speech alignment. Then, we utilize the speech-to-speech dataset to train the model to learn speech understanding and speech synthesis capabilities. At this stage, only the parameters of the audio decoder are trainable.
3.4 Nexus-O-audio testbed
Our Nexus-O-audio testbed comprises audio samples collected from real-world scenarios, including video meetings, live streaming, voice-based customer service QA interactions, and real-life conversations. We apply an off-the-shelf ASR model to convert each speech into corresponding text as our testbed. More information is provided in Figure 3. In addition to the ASR task, we plan to introduce Audio QA, Audio Visual QA tasks, and tri-modal tasks, which are currently under development.
4 Evaluation
We systematically evaluate our model performance from two perspectives: (i) downstream tasks alignment analysis in Section 4.1 and (ii) representation space alignment analysis in Section 4.2.
4.1 Multimodal Tasks Alignment Analysis
In this section, we evaluate the performance of Nexus-O on the following four tasks: vision understanding, audio English tasks, ASR, and S2TT. These tasks cover a wide range of modalities and capabilities, providing a comprehensive assessment of the versatility and effectiveness of MLLMs across various domains.
4.1.1 Vision-Language Evaluation
We first quantitatively evaluate Nexus-O on various vision understanding tasks, including the following benchmarks: (1) HallusionBench (Hal) (Guan et al., 2023): evaluating the hallucination and visual illusion. (2) MathVista (MathV) (Lu et al., 2023): evaluating mathematical reasoning in visual contexts. (3) OCRBench (OCR) (Fu et al., 2024d): evaluating Optical Character Recognition (OCR) capabilities on text-intensive images. (4) Video-MME (Fu et al., 2024b): evaluating video QA reasoning capabilities of MLLMs over 6 domains, diverse time range, etc. (5) MMMU (Yue et al., 2024): evaluating MLLMs on massive multi-discipline tasks demanding college-level subject knowledge and deliberate reasoning. (6) AI2D: AI2 Diagrams (AI2D) is a dataset of over 5000 grade school science diagrams. (7) MMVet (Yu et al., 2024): evaluating the integrated capabilities of MLLMs, including recognition, OCR, knowledge, language generation, spatial awareness, and math. (8) MME (Fu et al., 2024a): measuring both perception and cognition abilities on a total of 14 subtasks.
As illustrated in Table 2, we observe that Nexus-O achieves strong performance to current popular vision-language MLLMs on the MathV, OCR, and Hal benchmarks. Furthermore, when compared to Qwen2.5-VL, Nexus-O demonstrates superior performance, highlighting its competitive capabilities in vision understanding and reasoning tasks, indicating our model can effectively maintain vision-language alignment capability after auditory pretraining stage.
Table 2: Evaluation on Vision Understanding Benchmarks. The best two values are shown in bold and underlined. The blue row refers to the main competitors. In MMMU benchmark, * indicates that the measurement is developed over the validation set. We can observe that Nexus-O shows performance comparable to the leading open-source models and advanced closed-source counterparts.
Model LLM-size Video-MME MMMU MathV Hal AI2D OCR MMVet MME
Vision-Language Models
VILA-1.5 Vicuna-v1.5-13B 44.2 41.1 42.5 39.3 69.9 460.0 45.0 1718.2
LLaVA-Next Yi-34B 51.6 48.8 40.4 34.8 78.9 574.0 50.7 2006.5
CogVLM2 Llama3-Instruct-8B 50.5 42.6 38.6 41.3 73.4 757.0 57.8 1869.5
InternLM-Xcomposer2 InternLM2-7B 56.2 41.4 59.5 41.0 81.2 532.0 46.7 2220.4
Cambrian NousHermes2-Yi-34B 54.2 50.4 50.3 41.6 79.5 591.0 53.2 2049.9
InternVL-Chat-1.5 InternLM2-20B 57.1 46.8 54.7 47.4 80.6 720.0 55.4 2189.6
Ovis1.5 Gemma2-It-9B 58.1 49.7 65.6 48.2 84.5 752.0 53.8 2125.2
InternVL2 InternLM2.5-7B 61.5 51.2 58.3 45.0 83.6 794.0 54.3 2215.1
MiniCPM-V 2.6 Qwen2-7B 57.5 49.8 60.6 48.1 82.1 852.0 60.0 2268.7
Qwen2.5-VL Qwen2.5-7B 56.0 51.8* 61.1 71.7 80.7 877.0 - 2299.1
Omni-modal Models
VITA-1.5-Audio Qwen2-7B - 52.1 66.2 44.9 79.3 732.0 49.6 2352.0
EMova-8B LLaMA-3.1-8B - - 61.1 - 82.8 824.0 55.8 2205.0
Baichuan-Omni-1.5 - 58.2 47.3 51.9 47.8 - - 65.4 2186.9
Megrez-3B-Omni Megrez-3B - 51.8 62.0 50.1 82.0 - - 2315.0
Proprietary
GPT-4V - 50.4 59.3 48.2 39.3 71.4 678.0 49.0 1790.3
GPT-4o mini - 54.8 60.0 52.4 46.1 77.8 785.0 66.9 2003.4
Gemini 1.5 Pro 200B 59.1 60.6 57.7 45.6 79.1 754.0 64.0 2110.6
GPT-4o - 61.6 62.8 56.5 51.7 77.4 663.0 66.5 2328.7
Claude3.5 Sonnet 175B 62.2 65.9 61.6 49.9 80.2 788.0 66.0 1920.0
Our Model
Nexus-O Qwen2.5-VL-7B 57.0 53.2* 62.1 71.1 81.2 882.0 - 2315.5
4.1.2 Audio-Language Evaluation
Next, we evaluate the audio-language alignment on English QA, ASR, and S2TT tasks.
English QA task.
We evaluate the performance on the spoken QA benchmark: LLaMA Q. (Nachmani et al., 2024). As presented in Table 3, our model achieves top performance, higher than same-period competitor MiniCPM-o2.6-7B (highlighted in blue), demonstrating its competitive capabilities in this task.
Table 3: Evaluation on Audio English QA Benchmarks. The accuracy (%) of different models in English question answering on three sets. The parameter size is derived from the backbone LLMs. Our model achieves top accuracy in the LLaMA Q. benchmark, outperforming the same-period competitor MiniCPM-o2.6-7B.
Model Modality LLaMA Q.$\uparrow$
SpeechGPT-7B(Zhang et al., 2023) Audio&Text 21.60
Spectron-1B(Nachmani et al., 2023) Audio&Text 22.90
Moshi-7B(Défossez et al., 2024) Audio&Text 62.30
GLM-4-Voice-9B(Zeng et al., 2024) Audio&Text 64.70
MiniCPM-o2.6-7B Audio&Text 61.00
Mini-Omni-0.5B(Xie & Wu, 2024a) Audio&Text 22.00
Llama-Omni-8B(Fang et al., 2024b) Audio&Text 45.30
Our Models
Nexus-O Audio&Text 67.33
ASR task.
In the ASR task, we focus on both Mandarin (cn) and English (eng), evaluating performance on the following benchmarks: AIShell-2 (Du et al., 2018b), Librispeech (Panayotov et al., 2015b), and our real-scenario benchmark. As demonstrated in Table 4, our model achieves the best performance on the real scenario benchmarks, indicating the robustness of our model.
Table 4: Evaluation on ASR Benchmarks. Nexus-O has demonstrated strong performance in both real-life Mandarin and English ASR tasks. It outperforms specialized speech models, achieving better results in both languages. AIShell-2 is measured over all categories (Mac/iOS/Android), TC: test-clean, TO: test-other. We can observe that Nexus-O outperforms other models in real scenarios.
[…]
AIShell-2 Nexus-O-audio Librispeech TC Librispeech TO Nexus-O-audio
Speech LLMs
Qwen2-Audio-Instruct-7B 3.00/3.00/2.90 35.45 1.70 4.00 26.12
Omni-modal LLMs
Mini-Omini2-0.5B(Xie & Wu, 2024a) 117.71 9.80 4.70 54.69
VITA-1.5-7B(Fu et al., 2025) - 127.34 7.20 3.40 122.71
MinMo-7B(Chen et al., 2025b) 4.89/4.76/4.96 - 1.74 3.89 -
MiniCPM-o2.6-7B - 12.00 1.70 3.89 18.00
Our Models
Nexus-O 3.35 11.82 3.50 5.29 12.32
Speech-to-Text translation task.
As for the S2TT, we exploit CoVoST2 (Wang et al., 2020), which provides massive multilingual S2TT datasets. As shown in Table 5, Nexus-O demonstrates superior performance compared to specialised speech LLM, Qwen2-Audio-7B-Instruct, in both translation tasks. Nevertheless, our model still falls short of the performance of its contemporary competitor, MiniCPM-o2.6-7B. However, we pre-train our model on only 8 million audio samples, which is a relatively modest dataset at the pretraining stage.
Table 5: Evaluation on Speech-to-Text Translation Benchmarks. Nexus-O demonstrated superior performance compared to Qwen2-Audio-Instruct-7B in both translation tasks, however, it still falls short of the performance of its contemporary competitor. We conjecture that this performance gap may be due to the relatively modest scale of our pretraining speech dataset.
[…]
CoVoST2 CoVoST2
Speech LLMs
Qwen2-Audio-7B 24.40 45.20
Qwen2-Audio-Instruct-7B* 22.90 39.50
MinMo-7B * 25.95 46.68
Omni-modal LLMs
MiniCPM-o2.6-7B 27.20 48.20
Our Models
Nexus-O 23.17 40.21
4.2 Representational Alignment Analysis
In this section, we evaluate the alignment from the perspective of representation space to offer a deep evaluation beyond conventional downstream task alignment. We formalise MLLM within the framework of an unembedding-embedding architecture. In this framework, the unembedding stage is responsible for learning transformations between observations (e.g., text, vision, audio) and latent spaces through encoders, while the embedding stage captures the complex interactions among latent variables within the latent space of the hidden layers in MLLMs. Each stage serves distinct functions and yields representations with different properties (Park et al., 2024; Zhang et al., 2024d). Consequently, by focusing on each stage independently, we can have a systematical evaluation of model behaviours in representation spaces. To assess the representational alignment among vision-language modalities at each stage, we next measure the geometrical similarity among them via the kernel-alignment metrics. More information is provided in Appendix B.
Unembedding stage.
In the unembedding stage, we evaluate representational alignment using two language-vision datasets: ImageNet (Deng et al., 2009) and Wikipedia Caption (WIT) (Srinivasan et al., 2021). These datasets offer varying levels of fine granularity in language-vision alignment, enabling a comprehensive assessment of representational performance. To get the audio input for each language-vision pair, we apply the pretrained Text-to-Speech LLMs, cosyvoice-300M-SFT checkpoint, (Du et al., 2024). As for the baseline, we compare our model with MLLMs after SFT, including Qwen2.5-VL (Bai et al., 2025), LLaMA3.2-Vision (Meta., 2024), LLaVA family Liu et al. (2023b; a; 2024), Phi-3.5-Vision Abdin et al. (2024), and InternVL-8B (Chen et al., 2024c). The kernel-alignment metrics involve cka (Kornblith et al., 2019), cknna (Huh et al., 2024), svcca (Raghu et al., 2017), and $k$-nearest neighbours (knn)-based metrics: cycle knn, mutual knn, lcs knn, and edit knn. More information for those metrics can be found in (Huh et al., 2024).
As shown in Table 6, our tri-modal model incorporating audio inputs can generally outperform other models across both corpora. This demonstrates that the inclusion of the audio modality enhances representational alignment between the language and vision modalities. Additionally, Figure 5 provides a Principal Component Analysis (PCA) visualisation of the vector spaces for different modalities. The visualisation highlights that each modality exhibits distinct latent space geometries, and concatenating the vision and audio spaces results in a fusion of previously separated latent space structures.
Table 6: Kernel alignment analysis for unembedding space. The representation for each sample is the averaged token embeddings. The best two values are shown in bold and underlined, we can observe our tri-modal model with audio inputs generally outperforms others on both datasets.
Baseline cycle knn mutual knn lcs knn cka cknna svcca edit knn
ImageNet: concepts
Qwen2.5-VL-7B 0.66053 0.04278 1.47807 0.07987 0.02483 0.09168 0.00100
Llama3.2-Vision-Instruct-11B 0.08608 0.04205 1.52788 0.06079 0.01403 0.11651 0.00061
LLaVA1.5-7B 0.48673 0.04164 1.43367 0.03363 0.01213 0.13553 0.00088
LLaVA1.6-7B 0.57173 0.02077 0.81645 0.08024 0.01577 0.15240 0.00037
Phi3.5-Vision-Instruct 0.02761 0.01257 0.52355 0.08614 0.00714 0.12118 0.00019
InternVL2-8B 0.08175 0.01637 0.72495 0.09185 0.00062 0.12148 0.00044
Nexus-O 0.67839 0.03418 1.25284 0.08758 0.02822 0.11963 0.00112
vision+audio 0.66432 0.03935 1.39794 0.09706 0.02833 0.16667 0.00139
Wikipedia Caption: short descriptive sentences
Qwen2.5-VL-7B 0.52148 0.07005 2.09765 0.10194 0.04112 0.19668 0.00173
Llama3.2-Vision-Instruct-11B 0.31347 0.03623 1.29980 0.00968 0.00779 0.22120 0.00050
LLaVA1.5-7B 0.57714 0.02886 1.03808 0.04632 0.01586 0.25615 0.00046
LLaVA1.6-7B 0.57812 0.03935 1.36523 0.07933 0.03998 0.23114 0.00082
Phi3.5-Vision-Instruct 0.04980 0.03027 1.14843 0.01669 0.03890 0.18183 0.00066
InternVL2-8B 0.36914 0.04132 1.55761 0.04732 0.01658 0.21739 0.00093
Nexus-O 0.48828 0.06990 2.07714 0.08924 0.04101 0.20311 0.00167
vision+audio 0.54000 0.10636 2.79003 0.12403 0.06329 0.25624 0.00230
Figure 5: The unembedding space is visualized using Principal Component Analysis (PCA), with distinct colours representing clusters identified by k-means clustering. The visualisation reveals that, despite being trained within a unified model, the different modalities exhibit distinct latent space geometries.
Embedding stage.
Next, we focus on the vision-language alignment at the embedding stage. We restrict our attention to the Wikipedia-caption corpus (Srinivasan et al., 2021). We compare our model exclusively with Qwen2.5-VL-7B. For each sample, the representation is the averaged token embeddings or the last token embeddings. Table 7 illustrates the kernel-alignment scores at the last hidden layer, we can observe that incorporating audio into the vision modality improves the language-vision alignment of MLLMs, indicating the audio modality can provide complementary information to vision and language modalities.
Table 7: Kernel alignment analysis for embedding space. Some baselines (InternVL, Llava, etc.) do not support a single modality as input. Therefore, we do not evaluate them. We can observe that by incorporating audio into the vision modality, the vision-language leads to better alignment.
Baseline cycle knn mutual knn lcs knn cka cknna svcca edit knn
avg token embeds at last hidden layer
Qwen2.5-VL-7B 0.51660 0.08490 2.35839 0.20811 0.04076 0.46152 0.00201
Nexus-O 0.77929 0.13716 3.35058 0.20746 0.10547 0.47505 0.00431
vision+audio 0.79687 0.16107 3.76074 0.25915 0.12871 0.52091 0.00578
last token embed at last hidden layer
Qwen2.5-VL-7B 0.57421 0.09572 2.57714 0.21278 0.04942 0.34796 0.00226
Nexus-O 0.77343 0.10203 2.72851 0.10599 0.07068 0.39663 0.00306
vision+audio 0.90429 0.12000 3.12402 0.12787 0.08621 0.41922 0.00367
Additionally, we present a comparative analysis between Qwen2.5-VL-7B and our model across all hidden layers. As illustrated in Figure 6, the incorporation of the audio modality significantly enhances vision-language alignment, particularly in the intermediate hidden layers. Furthermore, we also provide a qualitative evaluation of language and vision alignment via PCA visualisation. In Figure 7, the observation reveals that the vision-audio fusion (depicted in the two images on the right) demonstrates a greater degree of overlap compared to the language-vision-audio separation (illustrated in the two images on the left). This suggests that the integration of audio enhances the alignment between language and vision modalities.
Figure 6: Averaged kernel-alignment score across different hidden layers, we can observe that audio modality (green bar) can result in better vision-language alignment at most hidden layers.
Figure 7: Qualitative evaluation for language-vision alignment at embedding hidden layers. We can observe that the vision-audio fusion (right two) has a larger overlap than language-vision-audio separation (left two), indicating incorporating audio leads to better language-vision alignment. we also present the PCA visualisation at all hidden layers in Figure 8 and 9.
5 Conclusion and Future Work
In this work, we propose an omni-perceive and -interactive tri-modal model, Nexus-O, that supports any combination of audio, image/video, and text inputs, and the Nexus-O-audio testbed. Unlike existing approaches that treat audio as an auxiliary, our model enables a joint understanding of these modalities. To facilitate the rapid deployment of the model in real-world scenarios, we investigate the synthesis of high-quality speech dialogue data augmented with actual environmental noise derived from public datasets. This approach substantially improves the performance and robustness of MLLMs in practical applications. To better understand the characteristics of diverse modalities throughout the integration process, we conducted an in-depth analysis from the perspective of representation space. As a cutting-edge contribution, we aim to establish a new standard for open-source multimodal foundation models and drive innovation in complex real-world applications.
In the future, we aim to enhance the capability of MLLMs to handle audio in a streaming manner, enabling real-time processing and generation of audio inputs and outputs for dynamic, time-sensitive applications. Additionally, we intend to integrate advanced Artificial Intelligence-Generated Content (AIGC) capabilities to allow the model to seamlessly generate high-quality, contextually coherent multimodal outputs, further broadening its applicability in creative and industrial domains.
Appendix A Experimental Details
The following tables describe the template we used for each pre-training task:
ASR <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|>Please help me transcribe my current speech into text. <|im_end|> <|im_start|>assistant [text] <|im_end|>
Speech-to-Text Chat <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> <|im_end|> <|im_start|>assistant [response] <|im_end|>
Speech-to-Text Translation <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> What is the English equivalent of this Chinese speech? <|im_end|> <|im_start|>assistant [response] <|im_end|>
Speech Synthesis <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>text<|audio_eos|> <|im_end|> <|im_start|>assistant [audio.wav] <|im_end|>
Speech-to-Speech Chat <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> <|im_end|> <|im_start|>assistant <|audio_bos|>audio.wav<|audio_eos|> <|im_end|>
Automatic Audio Caption <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> please describe the speech. <|im_end|> <|im_start|>assistant [text] <|im_end|>
Speech Emotion Recognition <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> what is the emotion in this speech? <|im_end|> <|im_start|>assistant [text] <|im_end|>
Appendix B Representational Alignment
Preliminaries.
Kernels, characterising the distance metrics between points in a representation space, are commonly used to assess vector space (Huh et al., 2024). Typically, the similarity between two kernels derived from different spaces indicates a higher degree of alignment between those spaces. This similarity can be quantified via kernel-alignment metrics, such as Centered Kernel Distance (cka) (Kornblith et al., 2019). Next, we formally define the target of our evaluation.
Given a data $x^{i}\in\mathcal{X}$ where $i=(1,\dots,N)$, each data can be represented as a corresponding tri-modal tuple $(x^{i}_{\text{language}},x^{i}_{\text{vision}},x^{i}_{\text{audio}})\in(\mathcal{X}_{\text{language}},\mathcal{X}_{\text{vision}},\mathcal{X}_{\text{audio}})$. For each data $x^{i}$, its vector representation can be mapped via feature function $f:\mathcal{X}\rightarrow\mathbb{R}^{n}$. At the unembedding stage, $f$ for different modalities are the corresponding encoders (i.e., $f_{\text{language}}$, $f_{\text{vision}}$, and $f_{\text{audio}}$). At the embedding stage, $f$ is the unified LLMs with different modal inputs, denoted by $f_{\text{LLM}}$.
For each modality, a kernel, which measures the distance/similarity between datapoints, can be denoted by $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. The measurement of a single modal space can be described as:
$K(\mathcal{X})=K(x^{i},x^{j})=\langle f(x^{i}),f(x^{j})\rangle$
where $\langle$,$\rangle$ , ⟩ denotes the inner product operation, $x^{i}$ and $x^{j}$ are different input points. In this experiment, we specifically assess the alignment between language-vision modalities by measuring the alignment between kernels through kernel-alignment metrics, $m$:
$m(K(\mathcal{X}_{\text{language}}),K(\mathcal{X}_{\text{vision}}))$
To further evaluate whether audio modality can assist vision-language representational alignment at each stage, we fuse vision and audio spaces and examine whether this fusion leads to better alignment to language space. In detail, for the unembedding stage, we concatenate $f_{\text{vision}}(x^{i}_{\text{vision}})$ and $f_{\text{audio}}(x^{i}_{\text{audio}})$ as the fused feature space. The kernel can be described as:
$K(\mathcal{X}_{\text{vision}+\text{audio}})=\langle[f(x_{\text{vision}}^{i});f(x_{\text{audio}}^{i})],[f(x_{\text{vision}}^{j});f(x_{\text{audio}}^{j})]\rangle$ (1)
For the embedding stage, we feed both vision and audio inputs into the model. The kernel can be described as:
$K(\mathcal{X}_{\text{vision}+\text{audio}})=\langle f_{\text{LLM}}(x^{i}_{\text{vision}},x^{i}_{\text{audio}}),f_{\text{LLM}}(x^{j}_{\text{vision}},x^{j}_{\text{audio}})\rangle$ (2)
The final kernel-alignment scores can be described as:
$m(K(\mathcal{X}_{\text{language}}),K(\mathcal{X}_{\text{vision}+\text{audio}}))$
Figure 8: Tri-modal alignment visualisation at different hidden layers (blue: language, orange: vision, green: audio).
Figure 9: Tri-modal alignment visualisation at different hidden layers (blue: language, green: vision+audio).