### v4

# NEXUS-O: An Omni-Perceptive And -Interactive Model for Language, Audio, And Vision

Abstract.
Human beings perceive the real world through a spectrum of sensory modalities, encompassing auditory, visual, and linguistic faculties. This work proposes an industry-level omni-modal large language model (LLM) pipeline that integrates auditory, visual, and linguistic modalities to overcome challenges such as limited tri-modal datasets, high computational costs, and complex feature alignments. Our pipeline consists of three main components: First, a modular, end-to-end framework enabling flexible configuration of various encoder–LLM–decoder architectures. Second, a lightweight training strategy that pre-trains audio-language alignment on the state-of-the-art vision-language model Qwen2.5-VL, thus avoiding the costly pre-training of vision-specific modalities. Third, an audio synthesis pipeline that generates high-quality audio-text data from diverse real-world scenarios, supporting applications such as Automatic Speech Recognition and Speech-to-Speech chat. To this end, we introduce an industry-level omni-modal LLM, NEXUS-O. Extensive experiments validate the efficacy of our pipeline, yielding the following key findings: (1) In the visual understanding task, NEXUS-O exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. (2) Within the English Spoken Question-Answering task, the model achieves better accuracy than the same-period competitor (i.e, MiniCPM-o2.6-7B) in the LLaMA Q. benchmark. (3) In our real-world ASR testset, NEXUS-O achieves outstanding performance, indicating its robustness in real scenarios. (4) In the Speech-to-Text Translation task, our model outperforms Qwen2-Audio-Instruct-7B. (5) In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0), NEXUS-O is comparable to its backbone vocoder on Seed-TTS benchmark. (6) An in-depth analysis of tri-modal alignment reveals that incorporating the audio modality enhances representational alignment between vision and language.
Omni-modalities, Mutimodal Large Language Models
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), Oct. 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754752isbn: 979-8-4007-2035-2/2025/10submissionid: MM-446ccs: Computing methodologies Natural language generation11footnotetext: $\dagger$ denotes project leader.11footnotetext: * denotes equal contribution.11footnotetext: $\ddagger$ denotes corresponding author.
1. Introduction
Since the mid-20th century, significant efforts have been devoted to designing systems of the capability to simulate human cognition, with the ultimate goal of developing Artificial General Intelligence (AGI). Notable algorithms include Rule-based systems, Neuro-symbolic AI systems, Neural networks systems, etc. Among them, neural networks systems have emerged as a cornerstone thanks to their remarkable generalization capabilities. In this context, various neural architectures have been proposed, such as CNN and Transformer. The discovery of neural scaling laws (Kaplan et al., 2020) has further advanced the field, paving the way for Transformer-based Large Language Models (LLMs), such as ChatGPT, which centre on language as the medium for perceiving and interacting with both human society and the physical world. By aligning multiple modalities around the language, Multimodal Large Language Models (MLLMs) have gained widespread recognition as a promising pathway toward AGI.
MLLMs, involving language, vision, and auditory modalities, can generally be divided into three categories: vision-language model (Wang et al., 2024), audio-language model (Chu et al., 2024a), and vision-language-audio model (Fu et al., 2025b). While alignment among three modalities offers a more comprehensive simulation of human cognitive processes than two-modality alignment, it presents significantly greater challenges. These challenges include limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. These limitations hinder the advancement of omni-perceptive and interactive AGI systems. To address these challenges, we propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline:
First, in the modularised framework, each modality (audio, image/video, and text) is processed using its own dedicated pretrained encoder. This design ensures that the inherent characteristics and fine-grained details unique to each modality are effectively preserved within their respective latent representations. Subsequently, these modality-specific representations are integrated and aligned within a LLM, and the resulting unified representations are subsequently supplied to decoupled decoders designed for each modality. This modular formalism design facilitates efficient and flexible configuration of various encoder–LLM–decoder architectures, thereby potentially contributing to the advancement of AGI by integrating components such as language models, world models, and embodied intelligence.
Second, most omni-modal models, such as Qwen2.5-omni(Xu et al., 2025), VITA1.5(Fu et al., 2024c), etc., necessitate a preliminary pre-training stage for vision alignment, understanding, and reasoning. This stage typically demands substantial computational resources due to the processing of long sequences of image tokens, which is consequently less accessible to smaller research institutions and companies. To mitigate this limitation, we only pre-train audio-language alignment over the current state-of-the-art vision-language model, Qwen2.5-VL (Bai et al., 2025) [^1] In this report, Qwen2.5-VL refers to the instruction version.. Due to the inherent alignment between audio and text modalities, this training strategy effectively integrates audio understanding and generation capabilities while preserving the model’s established vision comprehension and reasoning functions.
Third, high-quality audio training data are often less accessible in both academy and industry. To address this limitation, we propose an audio synthesis pipeline that generates high-quality audio-text data spanning a broad spectrum of real-world scenarios, such as live streaming and dialogue, thereby supporting various downstream tasks such as Automatic Speech Recognition (ASR) and Speech-to-Speech chat. By incorporating real-world audio data at the pretraining stage, the model can be fast deployed for real-world applications. Based on our audio synthesis pipeline, we introduce a real-life ASR testset, named as NEXUS-O-audio.
To this end, we introduce NEXUS-O222“Nexus”: connect the strongest models to form an even more powerful integrated model., an industry-level omni-modal LLM. Extensive experiments validate the efficacy of our pipeline, yielding the following key findings: (1) In the visual understanding task, NEXUS-O exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. (2) Within the English Spoken Question-Answering task, the model achieves better accuracy than the same-period competitor (i.e, MiniCPM-o2.6-7B) in the LLaMA Q. benchmark (Nachmani et al., 2024). (3) In our real-world ASR testset, NEXUS-O achieves outstanding performance, indicating its robustness in real scenarios. (4) In the Speech-to-Text Translation task, our model outperforms Qwen2-Audio-Instruct-7B. (5) In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0) NEXUS-O is comparable to its backbone vocoder on Seed-TTS benchmark (Anastassiou et al., 2024).
To further investigate the interplay among three modalities, we conducted an in-depth analysis of tri-modal alignment. Our findings indicate that (6) the inherent correspondence between audio and language can serve to facilitate and reinforce vision-language alignment, indicating the efficiency of our training strategy. The main contributions of this paper are summarised as follows:
  • We propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline.
  • Based on the proposed pipeline, we introduce NEXUS-O, an industry-level omni-modal LLM. It is designed to support any combination of audio, image/video, and text inputs, and it is capable of generating outputs in either audio or language modalities. The pretrained model is available at https://github.com/HiThink-Research/NEXUS-O
  • We systematically design the speech data synthesis pipeline to obtain high-quality, real-life speech datasets, covering various real-world scenarios, such as corporate meetings, live broadcasting scenarios, etc., to facilitate the rapid deployment of the model in real-world scenarios. Moreover, we introduce the NEXUS-O-audio testset to evaluate its robustness in real life ASR task.
  • We comprehensively evaluate the performance of NEXUS-O on different benchmarks and conduct an in-depth analysis of vision-language alignment by incorporating audio modality within the representation space. Experimental results indicate the merits of our proposed pipeline.
2. Related Work
In this section, we present the existing works of Multimodal Large Language Models (MLLMs) in Section 2.1, and analyse existing audio modality evaluation approaches in Section 2.2.
2.1. MLLMs
Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and reasoning over textual knowledge (Yang et al., 2024b). Building on these advancements, a bunch of studies (Dai et al., 2022) have extended the understanding and alignment capabilities of LLMs to visual knowledge, addressing challenges such as vision-language alignment and instruction-following, by introducing Multimodal Large Language Models (MLLMs). Notable models include Qwen-VL (Wang et al., 2024), InternVL (Chen et al., 2024b), etc. However, in complex visual reasoning tasks, such as video analysis, MLLMs typically struggle to perform well due to the heterogeneity among diverse modalities, where different modalities have distinct feature granularities, representations, and geometries in a shared latent space (Zhang et al., 2024). To further improve their performance in visual understanding tasks, recent works have expanded MLLMs by incorporating audio capabilities, such as LLaMA-Omni (Fang et al., 2024a), VITA-1.0 (Fu et al., 2024c), Mini-Omni2 (Xie and Wu, 2024a), Baichuan-Omni (Li et al., 2024a), and MinMo (Chen et al., 2025b). However, these models exhibit certain limitations: LLaMA-Omni and MinMo lacks visual capabilities, Baichuan-Omni does not have end-to-end Text To Speech (TTS) capabilities, and VITA-1.0 is constrained by the limited capability of its backbone model.
Recent models, such as VITA-1.5 (Fu et al., 2025b), MiniCPM-o2.6  [^3] https://github.com/OpenBMB/MiniCPM-o, Baichuan-omni-1.5444https://github.com/baichuan-inc/Baichuan-Omni-1.5, and OpenOmni (Luo et al., 2025), have been proposed and have achieved state-of-the-art performance. However, these models usually require a visual pretraining stage. While pretraining from scratch can enhance multimodal alignment, this strategy significantly increases training time and resource demands due to long sequence of image or video tokens and large pretraining vision data, which is less friendly with institutions with limited computation resources. To improve training efficiency, we start by pre-training over MLLM, specifically Qwen2.5-VL-7B. Owing to the natural alignment between audio and language modalities, the integration of the audio modality has minimal impact on the pretrained geometric and manifold structures underlying the vision-language alignment. As a result, the model’s established capabilities in vision comprehension and reasoning are effectively preserved.
2.2. Audio Modality Evaluation
With the rapid advancement of MLLMs, a wide range of multimodal benchmarks has been proposed and widely adopted for evaluation, ranging over vision (Fu et al., 2024b), audio (Fu et al., 2024a), omni (Li et al., 2024b), etc. In the speech domain, existing benchmarks, e.g., Fleurs (Conneau et al., 2023), Aishell2 (Du et al., 2018a), LibriSpeech (Pratap et al., 2020a), and Common Voice (Ardila et al., 2019), provide linguistically diverse datasets under varying acoustic conditions, facilitating Automatic Speech Recognition (ASR) evaluation. Additionally, AIR-Bench (Yang et al., 2024a) introduces a pioneering generative evaluation framework encompassing speech, natural sounds, and music, employing novel audio mixing strategies and GPT-4 for unified, objective, and reproducible assessments. However, the assessment of speech models should be closely aligned with real-world environments to ensure that the evaluation results faithfully represent the demands and requirements of practical and industrial applications. We observe that these existing speech benchmarks present three limitations. First, the small data scale of these benchmarks typically results in significant performance variance. Second, the benchmarks primarily focus on controlled scenarios, making it difficult to evaluate models in dynamic and complex real-world settings, e.g., meetings or live broadcast environments. Third, quantitative evaluation scores from the existing ASR benchmarks generally fail to align closely with the actual user interaction experience, thereby limiting their practical utility. To address these limitations, we propose NEXUS-O-audio testset, comprising 8.1k and 2.8k of Chinese and English ASR samples, spanning various application domains, such as corporate meetings and live broadcasting scenarios.
Figure 1. Modularised Architecture, which is designed to accept any combination of input modalities and generates output in either the language or audio modality, where the Auto-Regressive (AR) audio decoder takes a special start token embedding and the last language embedding as input to generate the hierarchical discrete audio codes in an auto-regressive manner. These codes are subsequently fed into a pretrained audio generator to produce the final audio output.
3. NEXUS-O
In this section, we introduce the architecture of NEXUS-O in Section 3.1, training data composition and audio data synthesis pipeline in Section 3.2, pre-training strategies in Section 3.3, respectively.
3.1. Modularised Architecture
Visual encoder.
Following Qwen2.5-VL (Bai et al., 2025), the vision encoder employs a re-designed Vision Transformer (ViT) architecture by incorporating M-RoPE and window attention to support native input resolutions while accelerating the computation of the entire visual encoder. During both training and inference, the height and width of the input images are resized to multiples of 28 before being fed into the ViT. The vision encoder processes images by splitting them into patches with a stride of 14, generating a set of image features.
Audio encoder/decoder.
To enable audio capabilities in the vision-language MLLM, we incorporate an audio encoder-decoder architecture. In this setup, the encoder is responsible for mapping the speech features into the semantic space of the MLLM, while the decoder transforms the semantic code back into speech, as shown in Figure 1. Specifically, the audio encoder comprises a pre-trained Whisper-large-v3 (Radford et al., 2023), and a two-layer MLP adapter. Specifically, the first layer of the MLP concatenates five consecutive speech feature vectors into a single vector, which is subsequently processed by the second MLP layer.
The auto-regressive audio decoder is a 6-layer decoder-only transformer designed to generate discrete speech codes, which are then passed to a pre-trained generator to produce the final waveforms. Specifically, we use the audio codebook from the FireFly GAN in Fish-Speech (Liao et al., 2024) as the audio codebook. Following MusicGen (Copet et al., 2023), the audio decoder predicts sequences of eight speech codes in a delayed manner. At each time step, the decoder averages the embeddings of the eight speech codes generated in the previous time step with the hidden state of the LLM at the current time step. The resulting averaged embedding is then fed into the audio transformer to predict the next eight speech tokens. Finally, the predicted speech tokens are input into the generator of the FireFly GAN to synthesise the final waveforms.
Language model.
NEXUS-O utilises the language model component of Qwen2.5-VL-7B (Bai et al., 2025) as the foundational language model.
3.2. Training Data
Figure 2. Audio dataset synthesis pipeline, the upper component represents the text-to-audio branch, while the lower component corresponds to the audio-to-text branch. Both components incorporate an equal proportion of in house, real-world samples.
Datasets composition.
As detailed in Table 1, the training data for NEXUS-O encompasses a wide range of domains within speech processing. These domains include (i) recognition tasks such as ASR and Speech Emotion Recognition (SER), (ii) reasoning tasks include Automatic Audio Caption and Speech Emotion Recognition, (iii) translation tasks like Speech-to-Text Translation (S2TT), generation tasks such as Speech Synthesis, and (iv) interactive tasks involving Speech-to-Speech Chat and Speech-to-Text Chat. The pretrained datasets include (1) public corpora (50%): Aishell1 (Bu et al., 2017), Wenetspeech (Zhang et al., 2022), Librispeech (Panayotov et al., 2015a), Kensho (O’Neill et al., 2021), PeopleSpeech (Galvez et al., 2021), MLS (Pratap et al., 2020b), and Gigast (Ye et al., 2023) and (2) real-life in-house data (50%).
Table 1 presents the statistical information for the audio training dataset. Unlike most Omni-MLLMs, which typically report duration in hours (Chen et al., 2025b), our summary is based on the number of audio samples. The total duration of the dataset is approximately 30k hours, this is considerably smaller than the durations employed in other works, such as MinMo (pretrained over 1.4 million hours) (Chen et al., 2025b) and Qwen2-Audio (pretrained over 520k hours) (Chu et al., 2024b).
Table 1. Datasets description.
Stage Task Description Num GPUs (H100)
1 Automatic Speech Recognition 5.6M 6
Speech-to-Text Translation 1.5M
Speech-to-Text Chat 0.2M
Automatic Audio Caption 0.14M
Speech Emotion Recognition 0.75M
2 Speech-to-Text Chat 0.23M 64
Automatic Speech Recognition 20k
Speech-to-Text Translation 20k
3 Speech Synthesis 40M 6
Speech-to-Speech chat 0.22M
Audio dataset synthesis pipeline.
We integrate both English and Chinese speech and textual datasets to generate paired speech-text data for ASR, speech synthesis (text-to-speech), Speech-to-Text Chat, and Speech-to-Speech Chat tasks. For the speech corpus (depicted at the bottom of Figure 2), we employ Whisper (Radford et al., 2022) to transcribe speech into text. For noisy samples in our in-house dataset, human annotators are engaged to ensure accurate transcription.
For the textural corpus (top in Figure 2), we employ the zero-shot in-context generation technique proposed by CosyVoice (Du et al., 2024a) to convert the text into natural-sounding speech. However, many conversational and instruction textural datasets are not ideal for speech synthesis due to the nature of the tasks, which do not facilitate effective speech interaction. To address this issue, we enhance the quality of our synthesised speech data by implementing the following pre-processing rules:
  1. (1)
    Length Filtering: texts exceeding a length threshold are excluded from the dataset. Specifically, texts longer than 200 Chinese characters or 200 English words are removed.
  2. (2)
    Non-text Element Filtering: texts containing an excessively high proportion of non-text elements are excluded. Specifically, texts with a non-text character ratio exceeding 0.2 are removed. Text characters in this context include punctuation marks (e.g., !), alphabetic characters ([a-z]), and Chinese characters.
  3. (3)
    Pattern Matching Filtering: texts that included URLs, file paths, or common LaTeX formulas were identified and removed using regular expressions.
For Speech-to-Text Chat and Speech-to-Speech Chat tasks, some texts are unsuitable for a conversational context. To address this limitation, we implement the following rules:
  1. (1)
    Excluding queries that are not appropriate for the Chat task. For example, questions involving metathetical questions are filtered out.
  2. (2)
    Rewriting the question in a conversational tone and generating a conversational response via Qwen2-72B (Chu et al., 2024a).
To enhance the diversity of voices, we randomly select a speaker from a pool of two thousand individuals for each corresponding language to synthesise the speech, ensuring that the voice aligns with the language of the text. Following the text-to-speech (TTS) process, an automated verification procedure is executed using the pretrained Whisper model, which checks whether the synthesised sample can be approximately reconverted to the original source sample, thereby ensuring the quality of the data synthesis.
Furthermore, we introduce the Nexus-audio testset, which comprises 8.1k Chinese and 2.8k English ASR samples. This testset spans a range of application domains, such as corporate meetings, live broadcasting scenarios, etc.
3.3. Training Strategies
Stage 1: audio alignment.
The first stage aims to facilitate the capacity of LLMs to comprehend and process input speech features. To accomplish this, we utilise 8 million bilingual speech-text pairs for alignment training, covering five training objectives, as illustrated in Table 1.
Stage 2: audio SFT
The second stage aims to enable the model to follow speech instructions by incorporating speech-based questions and text-based answers. In this stage, we freeze the visual and audio encoders while unfreezing the visual and audio adapters, and backbone LLM. This strategy is designed to enhance the capability of NEXUS-O to respond accurately to multimodal instructions, thereby improving its performance in understanding and processing both visual and auditory information.
By keeping the core encoder parameters fixed, we ensure that the foundational representations remain stable while allowing the adapters to fine-tune and specialize in the nuances of specific tasks or instructions. This strategic adjustment aims to refine the sensitivity of NEXUS-O to multimodal cues without disrupting the robustness of its underlying architecture.
Stage 3: audio output.
Different from the previous large speech model connected to the TTS module (Fu et al., 2024c), we exploit a decoder to generate speech end-to-end. First, we train NEXUS-O on the Text-to-Speech dataset to make the model learn text-to-speech alignment. Then, we utilize the speech-to-speech dataset to train the model to learn speech understanding and speech synthesis capabilities. At this stage, only the parameters of the audio decoder are trainable.
4. Evaluation
We systematically evaluate our model performance from two perspectives: (i) downstream tasks alignment analysis in Section 4.1 and (ii) representation space alignment analysis in Section 4.2.
4.1. Multimodal Tasks Alignment Analysis
In this section, we evaluate the performance of NEXUS-O on the following five tasks: vision understanding, Spoken English QA tasks, ASR, Speech-to-Text translation, and Text-to-Speech tasks. These tasks cover a wide range of modalities and capabilities, providing a comprehensive assessment of the versatility and effectiveness of MLLMs across various domains.
Vision-Language Evaluation
We conduct a comprehensive quantitative evaluation of NEXUS-Oon a range of vision understanding tasks using eight distinct benchmarks: (1) HallusionBench (Hal) (Guan et al., 2023): Evaluates the model’s susceptibility to hallucination and its handling of visual illusions. (2) MathVista (MathV) (Lu et al., 2023): Assesses mathematical reasoning within visual contexts. (3) OCRBench (OCR) (Fu et al., 2024d): Measures optical character recognition (OCR) capabilities on text-intensive images. (4) Video-MME (Fu et al., 2024b): Examines video question-answering reasoning abilities across six domains over diverse temporal ranges. (5) MMMU (Yue et al., 2024): Challenges the model with multi-disciplinary tasks that demand college-level subject knowledge and deliberate reasoning. (6) AI2D: A dataset comprising over 5,000 grade school science diagrams. (7) MMVet (Yu et al., 2024): Evaluates integrated capabilities including recognition, OCR, domain knowledge, language generation, spatial awareness, and mathematical reasoning. (8) MME (Fu et al., 2024a): Measures both perceptual and cognitive abilities across 14 subtasks.
Table 2. Evaluation on Vision Understanding Benchmarks. The best two values are shown in bold and underlined. The blue row refers to the main competitors. In MMMU benchmark, * indicates that the measurement is developed over the validation set. We can observe that NEXUS-O shows performance comparable to the leading open-source models and advanced closed-source counterparts.
Model LLM-size Video-MME MMMU MathV Hal AI2D OCR MMVet MME
Vision-Language Models
VILA-1.5(Lin et al., 2024) Vicuna-v1.5-13B 44.2 41.1 42.5 39.3 69.9 460.0 45.0 1718.2
LLaVA-Next(Liu et al., 2024) Yi-34B 51.6 48.8 40.4 34.8 78.9 574.0 50.7 2006.5
InternLM-Xcomposer2(Dong et al., 2024) InternLM2-7B 56.2 41.4 59.5 41.0 81.2 532.0 46.7 2220.4
Cambrian(Tong et al., 2024) NousHermes2-Yi-34B 54.2 50.4 50.3 41.6 79.5 591.0 53.2 2049.9
InternVL-Chat-1.5(Chen et al., 2024a) InternLM2-20B 57.1 46.8 54.7 47.4 80.6 720.0 55.4 2189.6
Ovis1.5(Lu et al., 2024) Gemma2-It-9B 58.1 49.7 65.6 48.2 84.5 752.0 53.8 2125.2
InternVL2(Chen et al., 2024b) InternLM2.5-7B 61.5 51.2 58.3 45.0 83.6 794.0 54.3 2215.1
MiniCPM-V 2.6(Yao et al., 2024) Qwen2-7B 57.5 49.8 60.6 48.1 82.1 852.0 60.0 2268.7
Qwen2.5-VL(Bai et al., 2025) Qwen2.5-7B 56.0 51.8* 61.1 71.7 80.7 877.0 - 2299.1
Omni-modal Models
VITA-1.5-Audio(Fu et al., 2025a) Qwen2-7B - 52.1 66.2 44.9 79.3 732.0 49.6 2352.0
EMova-8B(Chen et al., 2025c) LLaMA-3.1-8B - - 61.1 - 82.8 824.0 55.8 2205.0
Baichuan-Omni-1.5(Li et al., 2024a) - 58.2 47.3 51.9 47.8 - - 65.4 2186.9
Megrez-3B-Omni(Li et al., 2025) Megrez-3B - 51.8 62.0 50.1 82.0 - - 2315.0
Proprietary
GPT-4V - 50.4 59.3 48.2 39.3 71.4 678.0 49.0 1790.3
GPT-4o mini - 54.8 60.0 52.4 46.1 77.8 785.0 66.9 2003.4
Gemini 1.5 Pro 200B 59.1 60.6 57.7 45.6 79.1 754.0 64.0 2110.6
GPT-4o - 61.6 62.8 56.5 51.7 77.4 663.0 66.5 2328.7
Claude3.5 Sonnet 175B 62.2 65.9 61.6 49.9 80.2 788.0 66.0 1920.0
Our Model
NEXUS-O Qwen2.5-VL-7B 57.0 53.2* 62.1 71.1 81.2 882.0 - 2315.5
As illustrated in Table 2, we observe that NEXUS-O achieves strong performance to current popular vision-language MLLMs on the MathV, OCR, and Hal benchmarks. Furthermore, when compared to Qwen2.5-VL, NEXUS-O demonstrates superior performance, highlighting its competitive capabilities in vision understanding and reasoning tasks, indicating our model and training strategy can effectively maintain vision-language alignment capability by incorporating audio modality (Finding1).
Audio-Language Evaluation
Next, we evaluate the audio-language alignment on Spoken Question-Answering (SQA), Automatic Speech Recognition (ASR), Speech-to-Text translation (S2TT), and Text-to-Speech (TTS) tasks.
SQA task
We evaluate the performance on the English spoken QA benchmark: LLaMA Q. (Nachmani et al., 2024). As presented in Table 3, our model achieves top performance, higher than same-period competitor MiniCPM-o2.6-7B (highlighted in blue), demonstrating its competitive capabilities in this task.
Table 3. Evaluation on Audio English QA Benchmarks. The accuracy (%) of different models in English question answering on three sets. The parameter size is derived from the backbone LLMs. Our model achieves top accuracy in the LLaMA Q. benchmark, outperforming the same-period competitor MiniCPM-o2.6-7B.
[…]
SpeechGPT-7B(Zhang et al., 2023) Audio&Text 21.60
Spectron-1B(Nachmani et al., 2023) Audio&Text 22.90
Moshi-7B(Défossez et al., 2024) Audio&Text 62.30
GLM-4-Voice-9B(Zeng et al., 2024) Audio&Text 64.70
MiniCPM-o2.6-7B Audio&Text 61.00
Mini-Omni-0.5B(Xie and Wu, 2024b) Audio&Text 22.00
Llama-Omni-8B(Fang et al., 2024b) Audio&Text 45.30
Our Models
NEXUS-O Audio&Text 67.33
ASR task
In the ASR task, we focus on both Mandarin (CH) and English (EN), evaluating performance on the following benchmarks: AIShell-2 (Du et al., 2018b), Librispeech (Panayotov et al., 2015b), and our real-scenario benchmark. As demonstrated in Table 4, our model achieves the best performance on the real scenario ASR testset, NEXUS-O-audio, indicating the robustness of our model in real life (Finding2).
Table 4. Evaluation on ASR Benchmarks, where CER and WER refer to Character Error Rate and Word Error Rate, respectively. NEXUS-O has demonstrated strong performance in both real-life Mandarin and English ASR tasks. It outperforms specialized speech models, achieving better results in both languages. AIShell-2 is measured over all categories (Mac/iOS/Android), TC: test-clean, TO: test-other.
[…]
AIShell-2 NEXUS-O-audio Librispeech TC Librispeech TO NEXUS-O-audio
Speech LLMs
Qwen2-Audio-Instruct-7B(Chu et al., 2024b) 3.00/3.00/2.90 35.45 1.70 4.00 26.12
Omni-modal LLMs
Mini-Omini2-0.5B(Xie and Wu, 2024b) 117.71 9.80 4.70 54.69
VITA-1.5-7B(Fu et al., 2025b) - 127.34 7.20 3.40 122.71
MinMo-7B(Chen et al., 2025a) 4.89/4.76/4.96 - 1.74 3.89 -
MiniCPM-o2.6-7B - 12.00 1.70 3.89 18.00
Our Models
NEXUS-O 3.35 11.82 3.50 5.29 12.32
S2TT task
As for the S2TT, we exploit CoVoST2 (Wang et al., 2020), which provides massive multilingual S2TT datasets. As shown in Table 5, NEXUS-O demonstrates superior performance compared to specialised speech LLM, Qwen2-Audio-7B-Instruct, in both translation tasks. Nevertheless, our model still falls short of the performance of its contemporary competitor, MiniCPM-o2.6-7B. However, we pre-train our model on a small audio dataset (about 30k hours) at the pre-training stage (Finding3).
Table 5. Evaluation on Speech-to-Text Translation Benchmarks. NEXUS-O demonstrated superior performance compared to Qwen2-Audio-Instruct-7B in both translation tasks, however, it still falls short of the performance of its contemporary competitor due to the relatively modest scale of our pre-training speech dataset (about 30k hours).
[…]
CoVoST2 CoVoST2
Speech LLMs
Qwen2-Audio-7B(Chu et al., 2024b) 24.40 45.20
Qwen2-Audio-Instruct-7B(Chu et al., 2024b)* 22.90 39.50
MinMo-7B(Chen et al., 2025b)* 25.95 46.68
Omni-modal LLMs
MiniCPM-o2.6-7B 27.20 48.20
Our Models
NEXUS-O 23.17 40.21
TTS task
Finally, we evaluate NEXUS-O on Seed-TTS benchmark (Anastassiou et al., 2024) [^5] https://github.com/BytedanceSpeech/seed-tts-eval. As shown in Table 6, our experiments demonstrate that when employing pretrained vocoders (e.g., Fishspeech1.4 or CosyVoice2.0), NEXUS-O achieves performance comparable to its competitors (e.g., CosyVoice1.0 or CosyVoice2.0) (Finding4).
Table 6. Evaluation on Text-to-Speech (TTS) Benchmarks. NEXUS-O demonstrated lower performance in both Mandarin (CH) and English (EN) TTS tasks. Notice: MinMo-7B’s result is based on their manually filtered samples as mentioned in their report (Chen et al., 2025b), therefore, we do not compare with it.
Model CH EN
[…]
CosyVoice1.0(Du et al., 2024b) 8.64 3.59
CosyVoice2.0(Du et al., 2024c) 4.61 2.43
CosyVoice2.0-SFT(Du et al., 2024c) 2.06 3.19
MinMo-7B(Chen et al., 2025b) 2.48 2.90
NEXUS-O (Fishspeech1.4) 7.55 13.02
NEXUS-O (CosyVoice2.0) 4.53 4.11
Thus far, our evaluations demonstrate that incorporating the audio modality enables the model to perform audio understanding and generation without compromising its vision comprehension and reasoning capabilities. In the following section, we provide an in-depth evaluation of the vision-language alignment from the perspective of representation space.
4.2. Representational Alignment Analysis
Table 7. Kernel alignment analysis for unembedding space. The representation for each sample is the averaged token embeddings. vision+audio refers to the concatenation between audio and vision embeddings. The best two values are shown in bold and underlined, we can observe that (1) our model without audio inputs can lead to better vision-language alignment than the backbone model on 5 out of 7 metrics, and (2) our model with audio inputs generally outperforms others on 4 out of 7 metrics.
Baseline cycle knn mutual knn lcs knn cka cknna svcca edit knn
ImageNet: concepts
Qwen2.5-VL-7B 0.66053 0.04278 1.47807 0.07987 0.02483 0.09168 0.00100
Llama3.2-Vision-Instruct-11B 0.08608 0.04205 1.52788 0.06079 0.01403 0.11651 0.00061
LLaVA1.5-7B 0.48673 0.04164 1.43367 0.03363 0.01213 0.13553 0.00088
LLaVA1.6-7B 0.57173 0.02077 0.81645 0.08024 0.01577 0.15240 0.00037
Phi3.5-Vision-Instruct 0.02761 0.01257 0.52355 0.08614 0.00714 0.12118 0.00019
InternVL2-8B 0.08175 0.01637 0.72495 0.09185 0.00062 0.12148 0.00044
NEXUS-O 0.67839 0.03418 1.25284 0.08758 0.02822 0.11963 0.00112
vision+audio 0.66432 0.03935 1.39794 0.09706 0.02833 0.16667 0.00139
We formalise MLLM within the framework of an unembedding-embedding architecture. In this framework, the unembedding stage is responsible for learning transformations between observations (e.g., text, vision, audio) and latent spaces through encoders, while the embedding stage captures the complex interactions among latent variables within the latent space of the hidden layers in MLLMs. Each stage serves distinct functions and yields representations with different properties. Consequently, by focusing on each stage independently, we can have a systematical evaluation of model behaviours in representation spaces.
To assess the representational alignment among vision-language modalities at each stage, we first extract embeddings for each sample’s modalities. Our evaluation focuses on two aspects: (1) whether our model inherently promotes superior vision-language alignment, and (2) whether incorporating the audio modality into the vision modality further enhances this alignment.
Quantitative evaluation of vision-language alignment is performed using kernel-alignment metrics. The kernel of a space encapsulates its underlying geometrical structure (i.e., distance metric), and the similarity between the kernels of different spaces serves as an indicator of the alignment between these spaces—in our case, across different modalities. Those metrics involve cka (Kornblith et al., 2019), cknna (Huh et al., 2024), svcca (Raghu et al., 2017), and $k$-nearest neighbours (knn)-based metrics: cycle knn, mutual knn, lcs knn, and edit knn. More information for those metrics can be found in (Huh et al., 2024).
Figure 3. Averaged kernel-alignment score across different hidden layers. vision+audio: both visual and auditory modalities are concurrently fed into the model. The final fused representation is the last token in the sequence as the comprehensive summary of the integrated modalities. we can observe that incorporating audio modality (green bar) can result in better vision-language alignment at most hidden layers.
Unembedding stage
In the unembedding stage, we evaluate representational alignment using ImageNet (Deng et al., 2009), which offers a fine granularity in language-vision alignment, enabling a detailed assessment of representational performance. To get the audio input for each language-vision pair, we apply the pretrained Text-to-Speech LLMs, cosyvoice-300M-SFT checkpoint (Du et al., 2024a). As for the baseline, we compare our model with MLLMs after SFT, including Qwen2.5-VL (Bai et al., 2025), LLaMA3.2-Vision (Meta., 2024), LLaVA family (Liu et al., 2023), Phi-3.5-Vision (Abdin et al., 2024), and InternVL-8B (Chen et al., 2024b).
As shown in Table 7, we can observe that (1) our model without audio inputs can lead to better vision-language alignment than the backbone model, Qwen2.5-VL-7B, on 5 out of 7 metrics , and (2) by incorporating audio modalities, our model outperforms others on 4 out of 7 metrics (Finding5). These results indirectly demonstrate that our modular formalism design concept can yield an efficient tri-modal alignment at the encoding stage.
Embedding stage
Next, we focus on the vision-language alignment at the embedding stage. We restrict our attention to the Wikipedia-caption corpus (Srinivasan et al., 2021). We present a comparative analysis between Qwen2.5-VL-7B and our model across all hidden layers. We feed both image and its audio into the model and consider the last token embedding as the final fused representation. As illustrated in Figure 3, the incorporation of the audio modality significantly enhances vision-language alignment, particularly in the intermediate hidden layers (Finding6). This finding underscores the efficacy of our training strategy, whereby pretraining exclusively on audio data yields a robust tri-modal alignment.
5. Conclusion and Future Work
In this work, we propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline to overcome the current challenges, including limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. To this end, we propose an industry-level omni-perceive and -interactive tri-modal LLM, NEXUS-O.
Extensive experiments validate the usefulness of our proposed pipeline and NEXUS-O. In the visual understanding task, NEXUS-O exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0), NEXUS-O is comparable to its backbone vocoder on Seed-TTS benchmark. In our real-world ASR testset, NEXUS-O achieves outstanding performance, indicating its robustness in real scenarios. In the future, we intend to link NEXUS-O with vision generative models under our modularised framework to allow the model to seamlessly generate high-quality, contextually coherent visual outputs. Integrating the Artificial Intelligence-Generated Content (AIGC) capabilities into the current model can further broaden its applicability in creative and industrial domains, such as Embodied Intelligence.